# Scalable Stewardship of an LLM-Assisted Clinical Benchmark with Physician Oversight

### v2

@@ 2.4 Phase 2: Independent recomputation of reference labels
Phase 2 proposes revised reference labels by independently recomputing each score from first principles, without access to the original benchmark label or its derivation artifacts. This phase aims to (i) construct a set of high-confidence updated labels for evaluation and downstream use, and (ii) quantify disagreement between independently recomputed labels and the original labels to triage cases for clinician adjudication in Phase 3. We applied the Phase 2 relabeling pipeline to the full MedCalc-Bench test split (n=1,047n=1{,}047) and to a stratified random sample of the training split (n=5,183n=5{,}183, ${\sim}50\%$ of the original 10053-instance training set) stratified by question type $q$.
@@ 2.5 Phase 3: Targeted physician adjudication
We compare $\hat{y}^{\mathrm{original}}$ and $\hat{y}^{\mathrm{new}}$ to $y^{\star}$ using (i) an agreement indicator and (ii) symmetric mean absolute percentage error (sMAPE, [5]) for answerable numeric cases. For real-valued outputs, agreement is defined as relative deviation within $\pm 5\%$ (i.e., $|\hat{y}-y^{\star}|/|y^{\star}|\leq 0.05$). For integer-valued ordinal scores, agreement is defined as $|\hat{y}-y^{\star}|\leq 1$ when $y^{\star}<20$. For abstentions, agreement requires an exact match (both $y^{\star}$ and $\hat{y}$ equal $\mathsf{N/A}$). For $N$ instances, sMAPE is computed as:
@@
Post-adjudication label finalization and qualitative analysis.
@@ 2.6 Controlled Study of Label Quality’s Downstream Impact
We finetuned two identical copies of Qwen3-8B, an open-weights LLM with 4.6 million downloads from Hugging Face in August 2025,[31] using Group Relative Policy Optimization (GRPO) which is a modern RL algorithm widely used in LLM alignment.[34] The algorithm iteratively modifies a language model’s weights to increase the probability of generating completions that receive a higher reward. In this framework, the reference label ($y^{\star}$, surrogated in practice by $\hat{y}$) directly specifies a rule-based reward function; changing $\hat{y}$ is therefore functionally equivalent to changing the model’s implicit alignment objective.
@@
More technical details are given in Supplementary Appendix §D.4 and §E.
Both models were trained on 4,593 training instances and evaluated on a held-out test set of 887 instances. To test the sensitivity of evaluation outcomes to the choice of reference labels, the primary outcome measure was the accuracy of both models’ computed scores when graded against the updated test labels ($\hat{y}^{\mathrm{new}}$).
@@ 3 Results
3.1 Phase 1
Audit statistics of §2.3.
Running five independent copies of the auditor agent on MedCalc’s test set (n=1,047) surfaced a substantial number of suspicious labels: out of 1,047 instances, the $\hat{y}^{\mathrm{original}}$’s of 279 (26.6%) were flagged as Likely Error by the audit, i.e. marked as Suspicious by at least 4 out of 5 auditor reviews. Flagged issues were distributed broadly across the different questions $q$, covering 40 out of the total 55 distinct scoring questions in the test set.
@@ 3.1 Phase 1
A physician author (AG) independently selected and reviewed 7 Suspicious instances related to his clinical specialty. For each instance, the physician read both $(C,q,\hat{y}^{\mathrm{original}})$ and the auditor agents’ critique of $\hat{y}^{\mathrm{original}}$. They agreed with the error identification in 7/7 reviews. The sample size is modest, but it suggested the clinical utility of running a larger-scale MedCalc label recomputation in Phase 2.
3.2 Phase 2
@@
Using the supermajority-based labeling approach described in §2.4 yielded high-confidence labels for 887 out of 1,047 MedCalc test instances (85%) and for 4,593 out of 5,183 sampled train instances (89%). In these high-confidence labels, the abstention rate, i.e. how often the finalized label is $\mathsf{N/A}$, was 7.4% on the test set and 5.7% on the train set. We retain such $\mathsf{N/A}$ instances in the relabeled benchmark, as they correspond to clinically relevant “cannot be computed” cases rather than noise.
@@
Among the 887 test instances with high-confidence $\hat{y}^{\mathrm{new}}$, we find that 220 (24.8%) exceed the triage threshold (rel.err) defined in §2.4, plus another 66 (7.4%) are flagged as $\mathsf{N/A}$. These implies a total of 27.3% (on the 1,047) of the original test instances as likely mislabeled. This is still likely an underestimate, as rel.err is a conservative, symmetric lower bound on the underlying relative error with respect to the latent $y^{\star}$. Moreover, the 160 test instances for which the labeling agents could not reach a supermajority (and are therefore deferred) are plausibly hard cases where both $\hat{y}^{\mathrm{original}}$ and $\hat{y}^{\mathrm{new}}$ are unreliable, potentially pushing the true mislabel rate above 27.3%.
@@
The close correspondence between Phase 1’s audit-based error rate (26.6%) and Phase 2’s recomputation-based estimate (27.3%) provides a sensitivity check. These phases differ meaningfully in their setup: Phase 1 audits labels with full metadata visibility, while Phase 2 recomputes without the anchoring. Still, they arrive at similar error estimates. While not constituting direct validation, this consistency suggests that the detected quality issues are robust to variations in prompting and workflow design, not dependent on a particular agentic configuration.
3.3 Phase 3
@@
Table 1: Original MedCalc test labels ($\hat{y}^{\mathrm{original}}$) and our recomputed labels ($\hat{y}^{\mathrm{new}}$), compared against 50 physician-computed labels ($y^{\star}$). sMAPE, short for symmetric Mean Absolute Percentage Error, reports the average relative percentage deviation from physicians’ labels; it is computed on 34 / 33 instances where each $(\hat{y}^{\mathrm{original}},y^{\star})$ / $(\hat{y}^{\mathrm{new}},y^{\star})$ pair are both numerical. “Agreement” counts how many predicted labels in a group match the physician label up to a reasonable threshold. Square brackets are bootstrap $95\%$ confidence intervals.
Metric $(\hat{y}^{\mathrm{original}},y^{\star})$ $(\hat{y}^{\mathrm{new}},y^{\star})$
sMAPE (%) 72.7 [47.9, 99.1] 20.1 [8.0, 35.9]
Agreement 10/50 (20.0%) 37/50 (74.0%)
@@
Table 1 quantifies the prediction accuracy of $\hat{y}^{\mathrm{original}}$ and $\hat{y}^{\mathrm{new}}$ on the set of test instances with physician-computed labels, which we have treated as $y^{\star}$. Across 50 sampled instances with high $\textsf{rel.err}(\hat{y}^{\mathrm{original}},\hat{y}^{\mathrm{new}})$ values, $\hat{y}^{\mathrm{new}}$ agrees with $y^{\star}$ 37 times, compared to $\hat{y}^{\mathrm{original}}$’s 10 times. Beyond binary agreement, the new labels drive sMAPE down from 72.7% (95% CI, 47.9-99.1) to 20.1% (95% CI, 8.0-35.9), meaning that $\hat{y}^{\mathrm{new}}$ typically lies much closer to physicians’ numerical assessments in both absolute and relative terms. Although the sample size is modest, these results provide a targeted reality check that the relabeled $\hat{y}^{\mathrm{new}}$’s are more aligned with expert judgment on a substantial fraction of the contentious instances triaged by Phase 2.
Estimated Error on the Entire Test Set.
Phase 2 flagged 286 of the 1,047 test instances as likely mislabeled. Because obtaining physician ground truth for the entire dataset is operationally infeasible, we estimated the error rate across the full test set using a conservative assumption that strongly favors the original benchmark. Specifically, we treated the original labels for all 761 unflagged cases as perfectly correct (0% error). Even granting this advantage, the projected error for the original labels is approximately 19.8%, whereas the error for our recomputed labels is at most 9.2%. This indicates that the new labels’ accuracy advantage is not an artifact of sampling but is statistically robust, even when the original labels are given the benefit of the doubt on the majority of the dataset. Derivation details are given in Appendix §C.
Post-adjudication qualitative analysis.
Retrospective comparison against single-blind physician recomputations suggests that a substantial share of label disagreement is driven by context underspecification: many original $(C,q)$ instances do not contain enough contextual qualifiers to define a unique clinical “ground truth” score. Two patterns recur. First, the benchmark rarely specifies a scoring timepoint. Many notes describe multi-visit trajectories over weeks or months, but the question does not clarify whether the score should be computed at presentation, peak severity, discharge, or another reference time (e.g. rows 138, 334, 341, 509, 569, 739, 761, 785). Second, scoring rules can vary across time and jurisdiction, so the “official” aggregation rule is not uniquely describable by a static benchmark label. For example, Morphine Milligram Equivalent (MME) conversions for methadone depend on which guideline is assumed: pre-2023 CDC guidance used a tiered rule, i.e. a multiplier of 12 for methadone doses >60>60 mg/day,[7] whereas the current CDC guideline adopts a flatter multiplier (4.7) while cautioning that conversion becomes unreliable at higher doses;[9] still, today’s evidence suggest that some state-level guidelines reflect the older CDC schedule.[25, 41] Under such ambiguity, multiple defensible answers exist and a single numeric gold label does not cleanly measure an LLM’s generalizable clinical capability.
@@ 3.4 Effect of Label Choice on Downstream RL Training
Averaging the final ten evaluation steps, the copy of Qwen3-8B trained on our recomputed labels ($\hat{y}^{\mathrm{new}}$) achieves a final test accuracy of $71.4\%$, whereas the policy trained on the original benchmark labels ($\hat{y}^{\mathrm{original}}$) plateaus at $62.6\%$. This constitutes a substantial $8.7\%$ absolute variation attributable solely to the variation of the reward assignment. Post-hoc sensitivity analysis we performed over a $3\times 3$ hyperparameter sweep confirms that this performance difference is directionally salient across all settings. We present the dynamics of test accuracy during training in Figure 3.
To contextualize this result, we note that a $8.7\%$ performance difference could eliminate or double the gaps between the top three reported alignment results on MedCalc-Bench (Table 3 in Supplementary Appendix D.2). It is also noteworthy that while the label discrepancy between $\hat{y}^{\mathrm{original}}$ and $\hat{y}^{\mathrm{new}}$ is roughly $27\%$3.2), the downstream performance degradation is smaller ($8.7\%$), suggesting that the RL process displays some resilience to noisy rewards, though a systematic stewardship and cleanup is still crucial.
More conservatively, this $\pm 8.7\%$ swing can be read as an empirical measure of evaluation uncertainty induced by label choice: changing only the reference labels–while holding the data, model class, and optimization fixed–can move the headline conclusion by $8.7\%$. When the “ground truth” is itself produced by black-box pipelines, neither estimator can be assumed superior a priori; absent physicians’ adjudication, benchmark scores begin to resemble measurements made with a yardstick with a stochastic offset. Phase 3 (§3.3) is therefore essential for restoring the evaluative purpose of the benchmark by anchoring contentious cases to physician judgment, so that downstream RL training and evaluation reflect clinically defensible targets rather than artifacts of label construction.
Figure 3: Label Instantiation Changes Alignment Interpretation. Test accuracy dynamics for Qwen3-8B trained via GRPO using the recomputed reference labels (green) versus the original MedCalc-Bench labels (grey). Shaded bands indicate $\pm 1\sigma$ smoothed over a 10-step window. Improving the reward signal’s factual grounding effects a $8.7\%$ absolute gain in the final moving averages of model test accuracy ($71.4\%$ vs. $62.6\%$).
@@ 4 Discussion
In this work, we show that label fragility in a partially LLM-labeled clinical workflow benchmark can meaningfully change both evaluation conclusions and model alignment outcomes. Estimation error can arise from feature extraction mistakes, mismatched or outdated aggregation rules, or question underspecification. For safety-critical workflows, the practical implication is not that a benchmark must be “perfect” at release. Rather, it is that the benchmark should be maintained as a living document, with labels, assumptions, and scoring criteria revisited as clinical practice evolves and as model capabilities change.
@@
A binding constraint in operationalizing this kind of stewardship is the finitude of medical domain experts’ attention. GPU hours and API calls are elastic to scale; physician hours are not. Any practical stewardship protocol must fit into an already stretched clinical environment, and it should minimize disruption to routine care delivery. We therefore frame stewardship as a hybrid oversight system: automated audit and independent recomputation provide broad screening at low marginal cost, and clinician adjudication is reserved for a small set of high-disagreement cases. The goal is not to replace clinician judgment, but to concentrate it where it is most likely to change labels and improve clinical alignment.
@@
Our qualitative findings highlight a failure mode that is easy to overlook. Specifically, we observed note-question pairs that are clinically incompatible, omit required inputs, or leave key qualifiers implicit (for example, the scoring timepoint). Other scores depend on guideline version or jurisdiction, so there may not be a single “official” rule implied by the prompt. In these settings, forcing a numeric answer is unsafe. An explicit abstention option, and evaluation rules that credit abstention when the score is not computable, should be treated as a safety guardrail rather than an edge-case exception.
@@
Beyond MedCalc-Bench, we view benchmark stewardship as one concrete way for academia and the healthcare community to help shape the trajectory of the next generation of clinical LLMs. Frontier model development is increasingly concentrated in a small number of well-resourced AI labs; this concentration can reduce direct feedback from diverse real-world clinical settings, affecting relevance and accuracy. Public benchmarks therefore play a major role in setting norms: they define the state of the art and train the intuitions of engineers who optimize against shared test sets. As a recent example, Appendix §G analyzes Anthropic’s “Claude for Healthcare” release, presenting evidence that their reported results likely relied on the original NeurIPS v1.0 dataset, the version we found to contain substantial label noise, rather than a later revision. This highlights a key vulnerability: when public benchmarks contain uncorrected errors, they may inadvertently obscure the true capabilities of frontier models, turning accountability tools into sources of measurement ambiguity.
Limitations of this study include the fact that physician adjudication provides an essential reference, but it was conducted on a targeted subset with a modest sample size. Larger, multi-institution replications would better characterize variability across clinicians and settings, and would enable estimation of inter-rater reliability for subjective cases. Our automated pipelines also depend on a particular verifier design and prompting strategy, which may miss correlated failure modes. In addition, our recomputed labels cover a high-confidence subset, and the hardest instances are plausibly those most likely to be deferred.
Label quality matters not only for measurement but also for shaping the next generation of language models. Benchmarks are increasingly reused for model selection and post-training; in this role, the yardstick becomes a teacher. Our controlled RL experiment isolates this effect by holding the training pipeline fixed and changing only the reward labels, producing a substantial gap in downstream accuracy. More broadly, concerns about synthetic-data feedback loops have motivated recent work on “model collapse,” where recursively generated data can implode future models if incorporated indiscriminately.[10, 36]
@@
Taken together, these findings argue for treating clinical NLP benchmarks built with partial LLM inputs as a living document like Wikipedia: transparent versioning, documented assumptions, explicit representation of “unknown,” and hybrid oversight that focuses clinician effort where it matters most. This is a pragmatic path to preserving the public-good role of benchmarks, so that they remain trustworthy yardsticks for evaluation, and safer teachers when they are reused to shape the models we deploy.
Supplementary Appendix
Scalable Stewardship of an LLM-Assisted Clinical
Benchmark with Physician Oversight
Junze (Tony) Ye, BS1,*; Daniel Tawfik, MD, MS2; Alex J. Goodell, MD, MS3; Nikhil V. Kotha, MD4; Mark K. Buyyounouski, MD, MS4; Mohsen Bayati, PhD1,4,5,*
1 Department of Operations, Information and Technology, Stanford Graduate School of Business
2 Department of Pediatrics, Division of Critical Care Medicine, Stanford University School of Medicine
3 Department of Anesthesiology, Perioperative and Pain Medicine, Stanford University School of Medicine
4 Department of Radiation Oncology, Stanford University School of Medicine
5 Department of Electrical Engineering, Stanford University School of Engineering
*Corresponding authors. Email: {junze, bayati}@stanford.edu

### v3

@@ 2.4 Phase 2: Independent recomputation of reference labels
Phase 2 proposes revised reference labels by independently recomputing each score from first principles, without access to the original benchmark label or its derivation artifacts. This phase aims to (i) construct a set of high-confidence updated labels for evaluation and downstream use, and (ii) quantify disagreement between independently recomputed labels and the original labels to triage cases for clinician adjudication in Phase 3. We applied the Phase 2 relabeling pipeline to the full MedCalc-Bench test split (n=1,047n=1{,}047) and to a stratified random sample of the training split (n=5,183n=5{,}183, ${\sim}50\%$ of the original 10,053-instance training set), stratified by calculator question type $q$ so that each of the 55 calculators is represented in approximately the same proportion as in the full training set.
@@ 2.5 Phase 3: Targeted physician adjudication
We compare $\hat{y}^{\mathrm{original}}$ and $\hat{y}^{\mathrm{new}}$ to $y^{\star}$ using (i) an agreement indicator and (ii) symmetric mean absolute percentage error (sMAPE, (Armstrong, 1985)) for answerable numeric cases. For real-valued outputs, agreement is defined as relative deviation within $\pm 5\%$ (i.e., $|\hat{y}-y^{\star}|/|y^{\star}|\leq 0.05$). For integer-valued ordinal scores, agreement is defined as $|\hat{y}-y^{\star}|\leq 1$ when $y^{\star}<20$. For abstentions, agreement requires an exact match (both $y^{\star}$ and $\hat{y}$ equal $\mathsf{N/A}$). For $N$ instances, sMAPE is computed as:
@@
Post-hoc label finalization and qualitative analysis.
@@ 2.6 Controlled Study of Label Quality’s Downstream Impact
We finetuned two identical copies of Qwen3-8B, an open-weights LLM with 4.6 million downloads from Hugging Face in August 2025 (Qwen, 2025), using Group Relative Policy Optimization (GRPO), a modern RL algorithm widely used in LLM alignment since its introduction in the DeepSeek-R1 work (Shao et al., 2024; Guo et al., 2025). Our setting admits a simplified, on-policy formulation of GRPO; for the reader’s convenience, we provide a self-contained derivation in Supplementary Appendix §D.4. The key property for our experiment is that the reference label ($y^{\star}$, surrogated in practice by $\hat{y}$) directly specifies a rule-based reward function; changing $\hat{y}$ is therefore functionally equivalent to changing the model’s implicit alignment objective.
@@
Full experimental details, including hyperparameters, hardware, and the sampling and evaluation protocol, are given in Supplementary Appendix §E.
Both models were trained on 4,593 training instances. We evaluate the resulting checkpoints along three dimensions: (1) accuracy on the 50 physician-labeled test instances ($y^{\star}$) from Phase 3, which serves as our primary measure; (2) accuracy on the full 887-instance held-out test set, graded against $\hat{y}^{\mathrm{new}}$; and (3) accuracy on three additional medical evaluation sets (MedQA and two recent MedCalc-Bench revisions) to test whether training-label effects generalize to related tasks. The primary result is reported in §3.4; full details are in Supplementary Appendix §E.
@@ 3 Results
3.1 Phase 1: Automated audit of original labels
Running five independent copies of the auditor agent on MedCalc’s test set (n=1,047) flagged 279 (26.6%) instances as as Likely Error, i.e. marked as Suspicious by at least 4 out of 5 auditor reviews. Flagged issues were distributed broadly across the different questions $q$, covering 40 out of the total 55 distinct scoring questions in the test set.
@@ 3.1 Phase 1: Automated audit of original labels
A physician author (AG) independently selected and reviewed 7 Suspicious instances related to their clinical specialty. For each instance, the physician read both $(C,q,\hat{y}^{\mathrm{original}})$ and the auditor agents’ critique of $\hat{y}^{\mathrm{original}}$. They agreed with the error identification in 7/7 reviews, suggesting the clinical utility of running a larger-scale MedCalc label recomputation in Phase 2.
3.2 Phase 2: Independent recomputation of labels
@@
The supermajority-based labeling approach discussed above (see Methods §2.4 for full details) yielded high-confidence labels for 887 out of 1,047 MedCalc test instances (85%) and for 4,593 out of 5,183 sampled train instances (89%). In these high-confidence labels, the abstention rate, i.e. how often the finalized label is $\mathsf{N/A}$, was 7.4% on the test set and 5.7% on the train set.
@@
Among the 887 test instances with high-confidence $\hat{y}^{\mathrm{new}}$, 220 (24.8%) differ from the original label by more than the $\pm 5\%$ (formally defined in Methods Eq. (1)), and another 66 (7.4%) are flagged as $\mathsf{N/A}$. Together, these imply that at least 27.3% of the 1,047 original test instances are likely mislabeled. This is plausibly an underestimate: the 160 test instances for which the labeling agents could not reach a supermajority are deferred because they are hard cases where both $\hat{y}^{\mathrm{original}}$ and $\hat{y}^{\mathrm{new}}$ are unreliable, and the true mislabel rate could be substantially higher once these are accounted for.
@@
The close correspondence between Phase 1’s audit-based error rate (26.6%) and Phase 2’s recomputation-based estimate (27.3%) provides a sanity check. These phases differ meaningfully in their setup: Phase 1 audits labels with full metadata visibility, while Phase 2 recomputes without the anchoring. Still, they arrive at similar error estimates. While not constituting direct validation, this consistency suggests that the detected quality issues are robust to variations in prompting and workflow design.
3.3 Phase 3: Physician adjudication of disagreements
@@
Figure 3: Original MedCalc-Bench labels ($\hat{y}^{\mathrm{original}}$) and our recomputed labels ($\hat{y}^{\mathrm{new}}$), compared against 50 physician-computed labels ($y^{\star}$). (a) sMAPE (symmetric Mean Absolute Percentage Error) measures the average relative deviation from physicians’ labels; error bars show bootstrap 95% confidence intervals. (b) Agreement counts how many labels match the physician label up to a reasonable threshold ($\pm 5\%$ for numeric, exact match for abstentions).
@@
Figure 3 compares the accuracy of $\hat{y}^{\mathrm{original}}$ and $\hat{y}^{\mathrm{new}}$ on 50 test instances with physician-computed labels ($y^{\star}$). $\hat{y}^{\mathrm{new}}$ agrees with $y^{\star}$ 37 out of 50 times (74%; 95% CI, 60–84%), compared to $\hat{y}^{\mathrm{original}}$’s 10 out of 50 (20%; 95% CI, 11–33%). Beyond binary agreement, the recomputed labels drive sMAPE down from 72.7% (95% CI, 47.9–99.1) to 20.1% (95% CI, 8.0–35.9), meaning that $\hat{y}^{\mathrm{new}}$ typically lies much closer to physicians’ numerical assessments. Although the sample size is modest, these results provide a targeted reality check that the relabeled $\hat{y}^{\mathrm{new}}$’s are more aligned with expert judgment on the contentious instances triaged by Phase 2.
A conservative extrapolation to the full test set and a post-hoc qualitative analysis of physician comments are provided in Supplementary Appendix §C.
Evaluating today’s LLMs with our curated labels.
Figure 4 compares several frontier models in February 2026, evaluated against reference labels we curated in our three-phase pipeline. Figure 4(a) shows model accuracy on the 50-instance subset for which we have physician-computed labels ($y^{\star}$). Performance on this subset is expectedly lower than on the full test set, since these are the instances flagged in Phase 2 as the hardest and most contentious cases. Figure 4(b) reports accuracy on the full 887 instances for which our pipeline produced high-confidence recomputed labels. Because physician ground truth is unavailable for most of these instances, the comparison here uses $\hat{y}^{\mathrm{new}}$ as a proxy for $y^{\star}$, which is justified by the Phase 3 evidence that $\hat{y}^{\mathrm{new}}$ aligns substantially more closely with physician judgment than $\hat{y}^{\mathrm{original}}$ does (Figure 3). Accuracies in Figure 4(b) are therefore measured against a pseudo-ground-truth, but the larger sample size yields tighter confidence intervals.
We note that for the same 887 instances, the original MedCalc-Bench labels are also available as a reference. Grading the same model outputs against the original labels yields substantially lower accuracies for all four models ($65.7\%$, $69.4\%$, $62.5\%$, $67.8\%$, in the same order as Figure 4(b)). This drop is not a difference in the models, but a direct consequence of label errors in the original benchmark, providing further evidence that the original labels meaningfully underestimate the capability of today’s frontier LLMs on medical score computation. These lower accuracy figures are consistent with numbers reported by industry: in Anthropic’s “Claude for Healthcare” launch (Anthropic, 2026a), our independent reproduction analysis (Supplementary Appendix §G) suggests that these figures were likely obtained using the original v1.0 labels, further illustrating how label noise in the benchmark can understate model performance even in high-profile industry evaluations.
(a) Physician-produced labels ($N=50$)
(b) Full-pipeline labels ($N=887$)
Figure 4: Frontier LLM accuracy ($\pm 95\%$ CI) on MedCalc-Bench test split under our newly curated reference labels. Models are evaluated with a uniform prompt template and built-in tool use (web search + code execution) under high reasoning effort. (a) Accuracy against physician-computed labels for a subset of 50 test instances flagged for manual review in Phase 3. (b) Accuracy on all 887 test instances our stewardship pipeline relabeled, combining labels curated from Phase 2 & 3.
@@ 3.4 Effect of Label Choice on Downstream RL Training
To test whether label quality affects not only evaluation but also model training, we evaluate the two RL-trained Qwen3-8B checkpoints (one trained on $\hat{y}^{\mathrm{original}}$, the other on $\hat{y}^{\mathrm{new}}$; see Methods §2.6) on the 50 physician-labeled test instances from Phase 3. Both models start from the same base checkpoint and achieve identical accuracy before training (28%). Figure 5 shows how their performance on physician-labeled instances evolves over the course of training. As training progresses, a clear gap emerges: during the final phase of training (steps 205–300), the model trained on $\hat{y}^{\mathrm{new}}$ achieves a mean accuracy of 51.9% (95% CI, 50.5–53.3) versus 38.4% (95% CI, 36.7–39.9) for the model trained on $\hat{y}^{\mathrm{original}}$, a difference of approximately 13.5 percentage points attributable solely to the choice of training reward labels.
Figure 5: RL-trained model performance on 50 physician-labeled test instances ($y^{\star}$). Two identical copies of Qwen3-8B are trained via GRPO, differing only in whether training rewards are derived from the original labels ($\hat{y}^{\mathrm{original}}$, grey) or our recomputed labels ($\hat{y}^{\mathrm{new}}$, green). Each box summarizes the distribution of per-checkpoint accuracies within the indicated step range; individual checkpoints (evaluated every 5 steps) are overlaid as dots. (a) Test accuracy against physician labels. (b) Format accuracy (whether the model’s output is parseable as a valid score).
When evaluated on the larger 887-instance held-out test set (graded against $\hat{y}^{\mathrm{new}}$), the gap narrows to +8.7+8.7 percentage points ($71.4\%$ vs $62.6\%$; see Appendix §E and Figure 9 for the full training dynamics). The smaller gap on the full test set is expected: the 50 physician-labeled instances were selected precisely because they had the highest disagreement between $\hat{y}^{\mathrm{original}}$ and $\hat{y}^{\mathrm{new}}$, so the effect of label choice is concentrated there. On the remaining instances, where the two label sets largely agree, the choice of training reward matters less, and the test-wide average reflects a mixture of both regimes. We also evaluate both checkpoints on three additional medical evaluation sets (MedQA and two recent MedCalc-Bench revisions). The advantage of training on $\hat{y}^{\mathrm{new}}$ persists, though the effect is smaller (approximately 1 percentage point; $p=0.001$). Importantly, training on higher-quality labels does not degrade performance on any of these related tasks; the improvement on the primary task comes without a trade-off elsewhere (Appendix §E.2).
All recomputed labels, physician adjudication results, raw pipeline outputs, and code to reproduce the RL experiments are publicly available at https://github.com/junzeye/validate-medcalc-labels.
@@ 4 Discussion
In this work, we show that label quality in LLM-assisted benchmarks can meaningfully change both evaluation conclusions and model alignment outcomes. LLM assistance in benchmark construction is not a shortcut that can be avoided; it is increasingly a necessity. As models become more capable, benchmarks must grow harder to remain informative, and the expert labor required to produce ground-truth labels at scale, particularly in domains like medicine where that labor comes from clinicians with limited availability, makes fully human annotation impractical. LLM-assisted labeling is therefore likely to become more common, not less. The practical implication is not that a benchmark must be “perfect” at release, but that it should be maintained as a living document, with labels, assumptions, and scoring criteria revisited over time: even though models used for labeling continue to improve, they still hallucinate and make errors, and as the broader community engages with a benchmark, new issues surface that warrant correction. More capable auditing tools and broader community feedback together make ongoing stewardship both necessary and increasingly feasible.
@@
A binding constraint in operationalizing this kind of stewardship is the finitude of domain experts’ attention. GPU hours and API calls are elastic to scale; expert hours are not. Any practical stewardship protocol must fit into an already stretched professional environment and should minimize disruption to routine workflows. We therefore frame stewardship as a hybrid oversight system: automated audit and independent recomputation provide broad screening at low marginal cost, and expert adjudication is reserved for a small set of high-disagreement cases. The goal is not to replace expert judgment, but to concentrate it where it is most likely to change labels and improve alignment. While we demonstrate this approach in medicine, the same design principle applies to any domain where expert labels are expensive and benchmark accuracy is safety-relevant, including legal reasoning, materials science, and financial risk assessment.
@@
Our qualitative analysis (Supplementary Appendix §C) reveals that a substantial share of label disagreement stems from context underspecification: benchmark instances that omit required inputs or leave key qualifiers implicit. In these settings, forcing a numeric answer is unsafe. Explicit abstention options, and evaluation rules that credit abstention when the answer is not computable, should be treated as a safety guardrail rather than an edge-case exception.
@@
Beyond MedCalc-Bench, we view benchmark stewardship as one concrete way for academia and the broader AI community to help shape the trajectory of the next generation of LLMs. Frontier model development is increasingly concentrated in a small number of well-resourced AI labs; this concentration can reduce direct feedback from diverse real-world settings, affecting relevance and accuracy. Public benchmarks therefore play a major role in setting norms: they define the state of the art and train the intuitions of engineers who optimize against shared test sets. When public benchmarks contain uncorrected errors, they may inadvertently obscure the true capabilities of frontier models, turning accountability tools into sources of measurement ambiguity.
Label quality matters not only for measurement but also for shaping the next generation of language models. Benchmarks are increasingly reused for model selection and post-training; in this role, the yardstick becomes a teacher. Our controlled RL experiment isolates this effect by holding the training pipeline fixed and changing only the reward labels, producing a 13.5 percentage-point gap in accuracy on physician-labeled instances. Importantly, training on higher-quality labels does not come at the cost of performance on related tasks: the advantage persists, albeit at a smaller magnitude, across additional medical evaluation sets. This finding connects to a broader concern in the AI community: when LLM-generated outputs are fed back into the training pipeline, whether as benchmark labels, synthetic training data, or reward signals, errors can compound across generations. This feedback loop has been studied under the heading of “model collapse,” where recursively generated data can degrade future models if incorporated without adequate quality controls (Gerstgrasser et al., 2024; Shumailov et al., 2024). Our results provide concrete evidence of a related mechanism in the benchmark setting: LLM-generated labels that are treated as static ground truth can silently bias the models trained against them.
Several limitations should be noted when interpreting these results. First, our physician ground truth comprises 50 instances, which is sufficient to detect the large effects we report but lacks statistical power to detect smaller differences. Second, we treat physician labels as ground truth, but clinicians themselves can make calculation errors or disagree on ambiguous cases. Ideally, each instance would be independently labeled by multiple clinicians from different institutions, and their responses aggregated to reduce noise in the reference standard; our single-physician design does not capture this inter-rater variability (Supplementary Appendix §D.3 discusses the implications of treating ground truth as a point estimate versus a distribution). Third, while our five-agent consensus mechanism in Phases 1 and 2 reduces stochastic noise, the agents are not fully independent: they share the same underlying model and prompting strategy, so systematic biases (e.g., a consistent misinterpretation of a scoring rule) would not be caught by majority voting. Finally, the RL experiment uses a single 8B-parameter model; larger models may exhibit different generalization behavior, and testing this is an important direction for future work.
@@
Taken together, these findings argue for treating LLM-assisted benchmarks as living documents: transparent versioning, documented assumptions, explicit representation of “unknown,” and hybrid oversight that focuses expert effort where it matters most. This is a pragmatic path to preserving the public-good role of benchmarks, so that they remain trustworthy yardsticks for evaluation, and safer teachers when they are reused to shape the models we deploy.
Acknowledgments
Some of the computing for this project was performed on the Stanford Marlowe GPU cluster (Kapfer et al., 2025). We thank Stanford University and Stanford Data Science for providing the computational resources at a subsidized rate that made these research results possible.